OpenStudy (anonymous):

verify this identity (a+b+c)^2= a^2+b^2+c^2+2ab+2bc+2ca

Parth (parthkohli):

It's always true.

OpenStudy (arbershabani97):

just multiply (a+b+c)*(a+b+c)= and just start solving and you get there, it's very simple

random231 (random231):

@arbershabani97 is right!

Parth (parthkohli):

\[(\overline{a + b} + c)^2 = (a + b)^2 + c^2 + 2(a + b)(c)\]

random231 (random231):

same thing lol :P

OpenStudy (anonymous):

\[(a+b+c)^{2}=a(a+b+c)+b(a+b+c)+c(a+b+c)\] Then distribute each one again and simplify :)

OpenStudy (amistre64):

is latex coding down? or just my end?

Parth (parthkohli):

Just you.

OpenStudy (amistre64):

refresh helped ....

random231 (random231):

ahem staph notifications lol :P